# Quantum automata and languages of finite index

### v2

@@ 2.3 Closure properties of semialgebraic sets
The next operation will be used. Given a bijection
@@
and a matrix $M\in\mathbb{R}^{n\times n}$ denote by $\pi(M)$ the matrix $\pi(M)_{i,j}=M_{\pi(i,j)}$. Extend this operation to subsets of matrices ${\cal A}$. Write $\pi({\cal A})$ to denote the set of matrices $\pi(M)$ for all $M\in{\cal A}$.
@@
Now we recall some closure properties of the class of effective semialgebraic matrices investigated in [7]. Let us first consider the operation (6). We adopt the following conventions: we write $\exists^{n}X$ to mean that $X$ is a vector of $n$ bound variables. Furthermore, a vector of $n\times n$ variables can be interpreted as an $n\times n$ matrix of variables. Consider two semialgebraic sets of matrices, say ${\cal A}_{1}$ and ${\cal A}_{2}$, respectively defined by first-order formulas $\Phi_{1}(X_{1})$ and $\Phi_{2}(X_{2})$ where $X_{1}$ and $X_{2}$ are two families of $n^{2}$ free variables viewed as two $n\times n$ matrices of variables. Then ${\cal A}_{1}{\cal A}_{2}$ is defined by the following formula where $X$ is a family of $n^{2}$ free variables viewed as an $n\times n$ matrix
$\exists^{n\times n}X_{1}\;\exists^{n\times n}X_{2}:X=X_{1}X_{2}\wedge\Phi_{1}(X_{1})\wedge\Phi_{2}(X_{2})$
and $X=X_{1}X_{2}$ is an abbreviation for the predicate defining $X$. This shows that the class of effective semialgebraic sets of matrices is closed with respect to (6). The verification of the properties stated below is done similarly.
@@
Let ${\cal A}_{1}\subseteq\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and ${\cal A}_{2}\subseteq\mathbb{R}^{n\times n}$ be semialgebraic (resp., effective semialgebraic). Then ${\cal A}_{1}\diamond{\cal A}_{2}$ is are semialgebraic (resp., effective semialgebraic).
@@ 2.4 Algebraic Irreducible sets
Let $\cal A$ be a subset of $O_{nk}$ with ${\cal A}\in({\cal H})$ and let $\Psi$ be a regular map
$\Psi:O_{nk}\;\;\displaystyle\mathop{\longrightarrow}\;\;O_{n}.$
If, for every $i=1,\ldots,k$, $\Psi({\cal A}_{i})$ contains $I$, then $\overline{\Psi({\cal A})^{*}}\in({\cal H})$. In particular $\Psi({\cal A)}$ is effective algebraic.
@@
Finally, let us define recursively the family of sets of $O_{n}$
$\{H_{i}\}_{i\geq 1},$
where $H_{0}=\{I\}$ and, for every $i\geq 1$,
@@
which, by the irreducibility of all of its terms, terminates, i.e., there exists some $m\in\mathbb{N}$ such that (cf Remark 1)
@@
Finally let ${\cal X}=\bigcup_{i\geq 1}\ H_{i}$. By (13), $\overline{\cal X}=H_{1}\cup\cdots\cup H_{m}$. On the other hand, since, for every $i\geq 1$, $H_{i}=\overline{(\Psi({\cal A}_{1})\Psi({\cal A}_{2}))^{i}}$, then it is readily checked that
$\overline{\cal X}=H_{1}\cup\cdots\cup H_{m}=\overline{\Psi({\cal A})^{*}}.$
@@ 3 Effectiveness issues
Equivalently we prove the inclusion $L\subseteq|{\cal Q}_{\leq}|$. We run in parallel two semialgorithms. The first one verifies the noninclusion by enumerating the words $w\in L$ and testing if $||s\varphi(w)P||>\lambda$ holds. The second semialgorithm considers an effectively computable polynomial $\phi(X)$ which defines ${\mathbf{C}l}(\varphi(L))$ and verifies whether the sentence
$\Psi\ \equiv\ \forall X:\phi(X)\Rightarrow sXP\leq\lambda$
holds, which can be achieved by Tarski Seidenberg elimination result. If the inclusion $L\subseteq|{\cal Q}_{\leq}|$ holds then the first semialgorithm cannot answer “yes” and the second semialgorithm will answer “yes”. Now assume that the inclusion does not hold. Then the first semialgorithm will eventually answer “yes”, while the second semialgorithm cannot answer “yes” since this would imply that $\Psi$ holds and thus
$\forall X:X\in{\mathbf{C}l}(\varphi(L))\Rightarrow||sXP||\leq\lambda,$
a contradiction. ∎∎
@@ 4 Preliminaries on context-free languages
Given a monoid $M,$ a subset of $M$ is said to be regular if it is finite or it is obtained from finite sets by using finitely many times the regular operations, i.e., the union of two sets, the product ()($\cdot$)( ⋅ ) of two sets, and the star $(^{*})$, this last operation associating with a set $\cal X$ the submonoid ${\cal X}^{*}$ generated by $\cal X$. It is worth noticing that the concept of regular set can be equivalently given in terms of $M$-automaton. We will recall this equivalence in the appropriate points (cf Lemma 6, Proposition 10).
A context-free grammar $G$ is a quadruple $\langle V,\Sigma,P,S\rangle$ where $\Sigma$ is the alphabet of terminal symbols, $V$ is the set of variables or nonterminal symbols, $P$ is the set of rules or productions, and $S$ is the axiom of the grammar. A word over the alphabet $\Sigma$ is called terminal. As usual, the nonterminal symbols are denoted by uppercase letters as, for instance, $A$, $B$. An arbitrary rule of the grammar is written as $A\rightarrow\alpha$, where $\alpha$ is a finite sequence of variables and terminals. The derivation relation of $G$ is denoted by $\mathop{\Rightarrow}^{*}$. The language generated by $G$ is the subset of terminal words $\{w\in\Sigma^{*}:S\displaystyle\mathop{\Rightarrow}^{*}w\}$ denoted by $L(G)$.
The family of context-free languages will be denoted by CFL.
Following [4], Sec. VII.5, we say that a context-free language $L$ is of index $k$, if there exist a context-free grammar $G$ with $L=L(G)$ and an integer $k\in\mathbb{N}$ such that the following property holds: for every word $w\in L(G)$, there exists a derivation $\delta=(S\displaystyle\mathop{\Rightarrow}^{*}w)$ such that the number of occurrences of variables in each sentential form of $\delta$ does not exceed $k$. The family of context-free languages of index $k$ will be denoted by $Ind(k)$ and ${\cal L}(IND_{FIN})$ will denote the family of all the languages of finite index, that is, $\bigcup_{k\in\mathbb{N}}Ind(k)$. One immediately sees that the most simple languages in such a family are those of $LIN$, the linear languages, i.e. those generated by grammars where the right hand side $\alpha$ of every rule $A\rightarrow\alpha$, contains at most one occurrence of nonterminal symbols, that is, $\alpha\in\Sigma^{*}\cup\Sigma^{*}V\Sigma^{*}$. A well-known result proven by Salomaa in [29] shows that Dyck languages cannot be generated by finite index grammars, thus implying ${\cal L}(IND_{FIN})\subset$CFL. The importance of the family ${\cal L}(IND_{FIN})$ is due to the fact that, by a well-known result by Baron and Kuich [2] (up to a technical restriction) the unambiguous context-free languages whose characteristic series in commutative variables are rational, are exactly those in ${\cal L}(IND_{FIN})$. For this reason, such languages are also called “quasi-rational” (see Section 4.2).
@@ 4.1 Structuring of CFL
$C_{A}\diamond L^{\prime}=\{uwv\mid(u,v)\in{C_{A}}\mbox{ and }w\in{L^{\prime}}\}$
As the proof of the main theorem proceeds by induction on the number of nonterminal symbols, we need to show how to recombine a grammar from simpler ones obtained by choosing an arbitrary non-axiom symbol as the new axiom and by canceling all the rules involving $S$. This is the reason for introducing the next notation.
@@
where $w_{1},\ldots,w_{\ell+1}\in\Sigma^{*}$ and $A_{1},A_{2},\ldots A_{\ell}\in V^{\prime}$. Next lemma easily follows from the previous definitions.
Lemma 3
With the notation of (16), the language $L^{\prime}(G)$ is the finite union of the languages
$w_{1}L_{A_{1}}w_{2}L_{A_{2}}\cdots w_{\ell}L_{A_{\ell}}w_{j_{\ell+1}}$
when $\beta$ ranges over all rules (15).
@@
With the previous notation $L$ is a finite union of languages of the form $C_{S}\diamond L^{\prime\prime}$ where
$L^{\prime\prime}=w_{1}L_{A_{1}}w_{2}L_{A_{2}}\cdots w_{\ell}L_{A_{\ell}}w_{\ell+1}$
Proof.
In order to prove the inclusion of the right- into left- hand side, it suffices to consider $w=\alpha u\beta,$ with $u\in L^{\prime}(G)$ and $(\alpha,\beta)\in C_{S}.$ One has $S\displaystyle\mathop{\Rightarrow}^{*}u$ and $S\displaystyle\mathop{\Rightarrow}^{*}\alpha S\beta$ so that $S\displaystyle\mathop{\Rightarrow}^{*}\alpha S\beta\mathop{\Rightarrow}^{*}\alpha u\beta,$ and thus $w\in L$.
Let us prove the opposite inclusion. A word $w\in L$ admits a derivation $S\displaystyle\mathop{\Rightarrow}^{*}w$. If the symbol $S$ does not occur in the derivation except at the start of the derivation, then $w\in L^{\prime}(G)$. Otherwise factor this derivation into $S\displaystyle\mathop{\Rightarrow}^{*}\alpha S\beta\mathop{\Rightarrow}^{*}w$ such that $S$ does not occur in the second part of the derivation except in the sentential form $\alpha S\beta$. Reorder the derivation $\alpha S\beta\mathop{\Rightarrow}^{*}w$ into $\alpha S\beta\mathop{\Rightarrow}^{*}\gamma S\delta\mathop{\Rightarrow}^{*}w$ so that $\gamma,\delta\in\Sigma^{*}$. This implies $w=\gamma u\delta$ for some $u\in L^{\prime}(G)$, which, by Lemma 3, completes the proof. ∎∎
@@ 4.2 Structuring of ${\cal L}(IND_{FIN})$
The following holds ([14], Theorem 4.2, cf also [4], Sec. VII.5, Theorem 5.2).
@@
For every $k\geq 1$, $Ind(k)=Qrt(k).$
4.2.1 Composition of grammars
For our next use, we will now describe how quasi-rational languages can be characterized in terms of composition of grammars.
For this purpose, we recall the construction of the grammar that generates the image $\theta(L)$ of a context-free language $L$ over the alphabet $\Sigma_{1}$ under a context-free substitution $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$, i.e. a $\cal F$-substitution, with $\cal F$=CFL.
We follow verbatim the argument of [4], Sec. II.2. Let ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ be the grammar that generates $L$ and let
${\cal G}_{2}=\{G_{x}\ :\ x\in\Sigma_{1}\}$
be the family of grammars where, for every $x\in\Sigma_{1}$,
$G_{x}=\langle V_{x},\Sigma_{2},P_{x},S_{x}\rangle$
is the grammar that generates $\theta(x)$. One may assume that $V_{1}\cap V_{x}=\emptyset$, for all $x\in\Sigma_{1}$. Let be the set
${\cal V}_{2}=\{S_{x}:x\in\Sigma_{1}\}$
and let the copy morphism
$\mbox{c}:(V_{1}\cup\Sigma_{1})^{*}\rightarrow(V_{1}\cup{\cal V}_{2})^{*},$ (18)
$G_{\mbox{c}}=\langle V_{1},\{S_{x}:x\in\Sigma_{1}\},P_{\mbox{c}},S\rangle,$
where $P_{\mbox{c}}=\{A\rightarrow\mbox{c}(\alpha):(A\rightarrow\alpha)\in P\}$. Finally, define
$H=\langle W,\Sigma_{2},Q,S\rangle$ (19)
as the grammar where
The grammar $H$ so obtained will be called the composition of ${\cal G}_{1}$ and ${\cal G}_{2}$ and denoted by
$G={\cal G}_{1}\circ{\cal G}_{2}.$ (20)
Remark 2
By a minor abuse of notation, up to the copy morphism (18), one may identify $\Sigma_{1}$ with ${\cal V}_{2}$. In particular ${V}_{1}\cap{\cal V}_{2}=\emptyset.$
The following lemma holds.
Lemma 4
$\theta(L)=L({\cal G})$.
A composition of $k$ families of grammars $({\cal G}_{1},\ldots,{\cal G}_{k}),$ with $k\geq 1$ will be the grammar
$G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k},$
obtained by iterating $(k-1)$ times the operation (20) of composition of grammars on the families of context-free grammars ${\cal G}_{1},\ldots,{\cal G}_{k}$ (if the operation is well defined). By (17), Theorem 4 can be formulated as follows.
Theorem 5
Let $L$ be an arbitrary language of $Qrt(k),\ k\geq 1$. Then $L$ is generated by a composition of $k$ families of linear grammars.
@@ 4.2.1 Composition of grammars
Observe that $\cal L$ is the image of the language $\{x^{n}:n\in\mathbb{N}\}$, $\Sigma_{1}=\{x\}$, under the substitution $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$, with $\theta(x)=\{a^{n}b^{n}:n\in\mathbb{N}\}$. Hence $\cal L$ is context-free of index $2$. Then $\cal L$ is generated by the composition $G={\cal G}_{1}\circ{\cal G}_{2}$ of the linear grammars ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ and ${\cal G}_{2}=\langle V_{2},\Sigma_{2},P_{2},\sigma\rangle$ where
@@ 5 Context-free languages and semialgebraic sets
Now we find convenient to recall (fixing some details of the proof) a property (cf [7], Prop. 7), obtained from the decomposition of Section 4.1. This property is relevant in this study since it states, for an arbitrary context-free language $L$, that the closure ${\mathbf{C}l}(\varphi(L))$ of the image of $L$, under the morphism (4) associated with a quantum automaton, is semialgebraic. To this purpose, the following construction is crucial. With an arbitrary nonterminal $A$ of $G$, associate the set $M_{A}$ defined as
@@
$(\varphi(u_{1})\oplus\varphi(u_{2})^{\scriptstyle T})(\varphi(v_{1})\oplus\varphi(v_{2})^{\scriptstyle T})=\varphi(u_{1}v_{1})\oplus\varphi(u_{2})^{\scriptstyle T}\varphi(v_{2})^{\scriptstyle T},$ (22)
@@
equals (22) for $\varphi(v_{2}u_{2})^{\scriptstyle T}=\varphi(u_{2})^{\scriptstyle T}\varphi(v_{2})^{\scriptstyle T}.$ Since $(\varepsilon,\varepsilon)\in M_{A}$, then $M_{A}$ is a monoid. The second statement comes from Theorems 2 and 3. ∎∎
@@
are algebraic. If $\mathop{\mathrm{Card}}\nolimits(V)=1$, i.e., the set of nonterminal symbols is reduced to $S$ then $L$ is reduced to $C_{S}\diamond L^{\prime}(G)$ and $L^{\prime}(G)$ is finite. We may further assume that there is a sole terminal rule $S\rightarrow w$, so that $L^{\prime}(G)=\{w\}$. By Theorem 1 one has [^3] By Proposition 1, if ${\mathbf{C}l}(M_{S})$ is algebraic, then $\{X\oplus Y^{\scriptstyle T}:X\oplus Y\in{\mathbf{C}l}(M_{S})\}$ is so
${\mathbf{C}l}(\varphi(L))=\{X\varphi(w)Y^{\scriptstyle T}\mid X\oplus Y\oplus\{\varphi(w)\}\in{\mathbf{C}l}(M_{S}\oplus\varphi(w))\}$ (26)
so that
@@
We recall that an intermediate class of languages in between the linear and the finite index ones, is that of metalinear (or ultralinear) languages ${\cal L}(Meta_{Lin})$. By definition, a language $L$ is in ${\cal L}(Meta_{Lin})$ if there exist a context-free grammar $G$ with $L=L(G)$ and an integer $k\in\mathbb{N}$ such that, in every derivation of $G$, the number of occurrences of variables in each sentential form does not exceed $k$. The following corollary holds.
@@
It is recursively decidable whether or not $L\cap|{\cal Q}_{>}|=\emptyset$ holds, if $L\in{\cal L}(Meta_{Lin})$ and $\cal Q$ is a rational quantum automaton.

### v3

@@ 2.3 Closure properties of semialgebraic sets
Given a bijection
@@
and a matrix $M\in\mathbb{R}^{n\times n}$ we define a new operation by denoting $\pi(M)$ as the matrix $\pi(M)_{i,j}=M_{\pi(i,j)}$. Extend this operation to subsets of matrices ${\cal A}$. Write $\pi({\cal A})$ to denote the set of matrices $\pi(M)$ for all $M\in{\cal A}$.
@@
Now we recall some closure properties of the class of semialgebraic sets of matrices (see [10], Sec. 2.4).
@@
Let ${\cal A}_{1}\subseteq\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and ${\cal A}_{2}\subseteq\mathbb{R}^{n\times n}$ be semialgebraic (resp., effective semialgebraic). Then ${\cal A}_{1}\diamond{\cal A}_{2}$ is semialgebraic (resp., effective semialgebraic).
Proposition 3
@@ 2.4 Algebraic Irreducible sets
Let $\cal A$ be a subset of $O_{m}$ with ${\cal A}\in({\cal H})$ and let $\Psi$ be a regular map
$\Psi:O_{m}\;\;\displaystyle\mathop{\longrightarrow}\;\;O_{n}.$
If, for every $i=1,\ldots,k$, $\Psi({\cal A}_{i})$ contains $I$, then $\overline{\Psi({\cal A})^{*}}\in({\cal H})$.
@@
Finally, let us define recursively the family of sets of $O_{n}$$\{H_{i}\}_{i\geq 1},$ where $H_{0}=\{I\}$ and, for every $i\geq 1$,
@@
which, by the irreducibility of all of its terms, terminates, i.e., there exists some $d\in\mathbb{N}$ such that
@@
Finally let ${\cal X}=\bigcup_{i\geq 1}\ H_{i}$. By (13), $\overline{\cal X}=H_{1}\cup\cdots\cup H_{d}$. On the other hand, since, for every $i\geq 1$, $H_{i}=\overline{(\Psi({\cal A}_{1})\Psi({\cal A}_{2}))^{i}}$, then it is readily checked that
$\overline{\cal X}=H_{1}\cup\cdots\cup H_{d}=\overline{\Psi({\cal A})^{*}}.$
@@ 3 Effectiveness issues
Equivalently we prove the inclusion $L\subseteq|{\cal Q}_{\leq}|$, which is equivalent to
$\forall\ w\in L\Rightarrow f(w)\ =\ ||s\varphi(w)P||^{2}\ \leq\ \lambda.$
By the continuity of the map $f$, the last is equivalent to
$\forall\ X\in{\mathbf{C}l}(\varphi(L))\ \Rightarrow\ ||sXP||^{2}\ \leq\ \lambda.$ (14)
By hypothesis, there exists an effectively constructible formula $\Phi(X)$ that defines ${\mathbf{C}l}(\varphi(L))$ as a semialgebraic set (see Definition 3). Hence (14) can be written as
$\Upsilon\;\equiv\;\forall\ X:\Phi(X)\ \Rightarrow\ ||sXP||^{2}\ \leq\ \lambda.$
This implies that $L\subseteq|{\cal Q}_{\leq}|$ if and only if $\Upsilon$ holds true. Finally, one verifies whether $\Upsilon$ holds, which can be achieved by applying the Tarski Seidenberg elimination result. ∎∎
@@ 4 Preliminaries on context-free languages
Given a monoid $M,$ a subset of $M$ is said to be regular if it is finite or it is obtained from finite sets by using finitely many times the regular operations, i.e., the union of two sets, the product ()($\cdot$)( ⋅ ) of two sets, and the star $(^{*})$, this last operation associating with a set $\cal X$ the submonoid ${\cal X}^{*}$ generated by $\cal X$. It is worth noticing that the concept of regular set can be equivalently given in terms of $M$-automaton. We will use this fact in the sequel (cf Lemma 4, Proposition 10).
A context-free grammar $G$ is a quadruple $\langle V,\Sigma,P,S\rangle$ where $\Sigma$ is the alphabet of terminal symbols, $V$ is the set of variables or nonterminal symbols, $P$ is the set of rules or productions, and $S$ is the axiom of the grammar. A word over the alphabet $\Sigma$ is called terminal. As usual, the nonterminal symbols are denoted by uppercase letters as, for instance, $A$, $B$. An arbitrary rule of the grammar is written as $A\rightarrow\alpha$, where $\alpha$ is a finite sequence of variables and terminals. The derivation relation of $G$ is denoted by $\mathop{\Rightarrow}^{*}$. The language generated by $G$ is the subset of terminal words $\{w\in\Sigma^{*}:S\displaystyle\mathop{\Rightarrow}^{*}w\}$ denoted by $L(G)$. The family of context-free languages is denoted by CFL.
Following [6], Sec. VII.5, we say that a context-free language $L$ is of index $k$, if there exist a context-free grammar $G$ with $L=L(G)$ and an integer $k\in\mathbb{N}$ such that the following property holds: for every word $w\in L(G)$, there exists some derivation $\delta=(S\displaystyle\mathop{\Rightarrow}^{*}w)$ such that the number of occurrences of variables in each sentential form of $\delta$ does not exceed $k$. The family of context-free languages of index $k$ will be denoted by $Ind(k)$ and ${\cal L}(IND_{FIN})$ will denote the family of all the languages of finite index, that is, $\bigcup_{k\in\mathbb{N}}Ind(k)$. One immediately sees that the most simple languages in such a family are those of $LIN$, the linear languages, i.e. those generated by grammars where the right hand side $\alpha$ of every rule $A\rightarrow\alpha$, contains at most one occurrence of nonterminal symbols, that is, $\alpha\in\Sigma^{*}\cup\Sigma^{*}V\Sigma^{*}$. A well-known result proven by Salomaa in [39] shows that Dyck languages cannot be generated by finite index grammars, thus implying ${\cal L}(IND_{FIN})\subset$CFL. The importance of the family ${\cal L}(IND_{FIN})$ is due to the fact that, by a well-known result by Baron and Kuich [3] (up to a technical restriction) the unambiguous context-free languages whose characteristic series in commutative variables are rational, are exactly those in ${\cal L}(IND_{FIN})$. For this reason, such languages are also called quasi-rational (see Section 4.2).
@@ 4.1 Structuring of CFL
$C_{A}\diamond L^{\prime}=\{uwv\mid(u,v)\in{C_{A}}\mbox{ and }w\in{L^{\prime}}\}.$ (15)
As the construction of the structuring mentioned above proceeds by induction on the number of nonterminal symbols, we need to show how to recombine a grammar from simpler ones obtained by choosing an arbitrary non-axiom symbol as the new axiom and by removing all the rules involving $S$. This is the reason for introducing the next notation.
@@
where $w_{1},\ldots,w_{\ell+1}\in\Sigma^{*}$ and $A_{1},A_{2},\ldots,A_{\ell}\in V^{\prime}$.
@@
([10], Prop. 20) The language $L$ is a finite union of languages of the form $C_{S}\diamond L^{\prime\prime}$ where
$L^{\prime\prime}=w_{1}L_{A_{1}}w_{2}L_{A_{2}}\cdots w_{\ell}L_{A_{\ell}}w_{\ell+1},$
when $\beta$ ranges over all rules (17), and with respect to the notation (18).
@@ 4.2 Structuring of ${\cal L}(IND_{FIN})$
The family $Qrt(k)$ is equivalently formulated in terms of the well known construct of composition of grammars (such construct is recalled in detail in the Appendix). The following holds.
@@
([18], Theorem 4.2, cf [6], Sec. VII.5, Theorem 5.2) Let $L$ be a context-free language and $k\geq 1$. Then $L\in Ind(k)$ if and only if $L\in Qrt(k)$. Moreover, $L\in Qrt(k)$ if and only if there exists a grammar $G$ with $L=L(G)$ such that $G$ is a composition ${\cal G}_{1}\circ\cdots\circ{\cal G}_{k}$ of $k$ families of linear grammars.
@@
Observe that $\cal L$ is the image of the language $\{x^{n}:n\in\mathbb{N}\}$, $\Sigma_{1}=\{x\}$, under the substitution $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$, with $\theta(x)=\{a^{n}b^{n}:n\in\mathbb{N}\}$. Hence $\cal L$ is a context-free language of index $2$. Then $\cal L$ is generated by the composition $G={\cal G}_{1}\circ{\cal G}_{2}$ of the linear grammars ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ and ${\cal G}_{2}=\langle V_{2},\Sigma_{2},P_{2},\sigma\rangle$ where
@@ 5 Context-free languages and semialgebraic sets
Now we find convenient to recall (fixing some details of the proof) a result (cf [10], Prop. 7), obtained from the decomposition of Section 4.1. Such a result is relevant in this study since it states, for an arbitrary context-free language $L$, that the closure ${\mathbf{C}l}(\varphi(L))$ of the image of $L$, under the morphism (4) associated with a quantum automaton, is semialgebraic. To this purpose, the following construction is crucial. With an arbitrary nonterminal $A$ of $G$, associate the set $M_{A}$ defined as
@@
$(\varphi(u_{1})\oplus\varphi(u_{2})^{\scriptstyle T})(\varphi(v_{1})\oplus\varphi(v_{2})^{\scriptstyle T})=\varphi(u_{1})\varphi(v_{1})\oplus\varphi(u_{2})^{\scriptstyle T}\varphi(v_{2})^{\scriptstyle T},$ (21)
@@
equals (21) for $\varphi(v_{2}u_{2})^{\scriptstyle T}=\varphi(u_{2})^{\scriptstyle T}\varphi(v_{2})^{\scriptstyle T}.$ Since $(\varepsilon,\varepsilon)\in C_{A}$, then $M_{A}$ is a monoid. The second statement comes from Theorems 2 and 3. ∎∎
@@
are semialgebraic. If $\mathop{\mathrm{Card}}\nolimits(V)=1$, i.e., the set of nonterminal symbols is reduced to $S$ then $L$ is reduced to $C_{S}\diamond L^{\prime}(G)$ and $L^{\prime}(G)$ is finite. We may further assume that there is a sole terminal rule $S\rightarrow w$, so that $L^{\prime}(G)=\{w\}$. By Theorem 1 one has [^3] By Proposition 1, if ${\mathbf{C}l}(M_{S})$ is algebraic, then $\{X\oplus Y^{\scriptstyle T}:X\oplus Y\in{\mathbf{C}l}(M_{S})\}$ is so
${\mathbf{C}l}(\varphi(L))=\{X\varphi(w)Y^{\scriptstyle T}\mid X\oplus Y\oplus\{\varphi(w)\}\in{\mathbf{C}l}(M_{S}\oplus\varphi(w))\}.$ (25)
On the other hand, one has
@@
An intermediate class of languages in between the linear and the finite index ones, is that of metalinear (or ultralinear) languages ${\cal L}(Meta_{Lin})$. By definition, a language $L$ is in ${\cal L}(Meta_{Lin})$ if there exist a context-free grammar $G$ with $L=L(G)$ and an integer $k\in\mathbb{N}$ such that, in every derivation of $G$, the number of occurrences of variables in each sentential form does not exceed $k$. The following corollary holds.
@@
If $L\in{\cal L}(Meta_{Lin})$ and $\cal Q$ is a rational quantum automaton, then it is recursively decidable whether or not $L\cap|{\cal Q}_{>}|=\emptyset$ holds,